# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

### v1

@@ Abstract
Automatic radiology report generation can alleviate the workload for physicians and minimize regional disparities in medical resources, therefore becoming an important topic in the medical image analysis field. It is a challenging task, as the computational model needs to mimic physicians to obtain information from multi-modal input data (i.e., medical images, clinical information, medical knowledge, etc.), and produce comprehensive and accurate reports. Recently, numerous works emerged to address this issue using deep learning-based methods, such as transformers, contrastive learning, and knowledge-base construction. This survey summarizes the key techniques developed in the most recent works and proposes a general workflow for deep learning-based report generation with five main components, including multi-modality data acquisition, data preparation, feature learning, feature fusion/interaction, and report generation. The state-of-the-art methods for each of these components are highlighted. Additionally, training strategies, public datasets, evaluation methods, current challenges, and future directions in this field are summarized. We have also conducted a quantitative comparison between different methods under the same experimental setting. This is the most up-to-date survey that focuses on multi-modality inputs and data fusion for radiology report generation. The aim is to provide comprehensive and rich information for researchers interested in automatic clinical report generation and medical image analysis, especially when using multimodal inputs, and assist them in developing new algorithms to advance the field.
@@ 1 Introduction
Medical images can offer detailed insights into bodies and help physicians screen, diagnose and monitor medical conditions without requiring invasive techniques (Beddiar et al., 2023; Liao et al., 2023). Radiologists summarize the information extracted from medical imaging into radiological reports for clinical decision-making. The manual generation of reports is however labour-intensive, time-consuming, and requires extensive expertise (Beddiar et al., 2023). Topol (2019) points out that the demand for medical image explanation greatly surpasses the current capacity of physicians in the United States. During an epidemic and with ageing populations, the situation can get worse. During the Covid-19 pandemic, for instance, in the UK, each radiologist was estimated to report as many as 100 images each day (Statistics, 2020). This makes it challenging for radiologists to provide high-quality reports within the scheduled time. The current demand extends patient waiting time and increases the risk of disease transmission (Beddiar et al., 2023) and compromises patient care. The development of automatic report generation techniques can help alleviate this problem.
Automatic high-quality report generation is challenging. It is intrinsically a multi-modality problem (Tu et al., 2024; Yan et al., 2023). In routine clinical practice, to produce clear, correct, concise, complete, consistent, and coherent reports, radiologists need to combine information from images with information from other modality data, such as clinical history and related clinical measures. Previously developed techniques mostly considered images as input, while for the past three years, multi-modality deep learning developed very rapidly. An increasing number of research papers endeavored to emulate physicians by leveraging multi-modal data for the generation of diagnostic reports, as shown in Figure 1.
@@
Most of the previous surveys on this topic (Kaur et al., 2022; Beddiar et al., 2023; Liao et al., 2023; Shamshad et al., 2023; Liu et al., 2023a) did not include non-imaging inputs. Messina et al. (2022) considered non-imaging inputs, but only involved 6 papers. Totally, the previous surveys included 40 to 66 papers for report generation, primarily focusing on articles published before 2022. This survey differs from previous ones in three main contributions: (1) we analyse an additional 22 papers that utilize non-image inputs, and focuses on the acquisition, analysis, and integration of multi-modal inputs. To the best of our knowledge, it is the first review to investigate state-of-the-art multi-modal data processing techniques for report generation; (2) we examine 89 papers published from 2021 to 2024 to provide a comprehensive study on novel techniques in automatic report generation; and (3) we propose a general workflow for report generation with a taxonomy of approaches employed, and we summarize training strategies, public datasets, and mainstream evaluation methods, as shown in Figure 2. The workflow includes 5 key components: inputs, data preparation, feature learning, feature fusion, and report generation. Table LABEL:tab:overall in Appendix A summarizes all works included in this survey.
Figure 2: The summarized workflow of automatic radiology report generation. The fundamental components and key techniques are included. The (x, y%) for each method represents the number and percentage of papers used that technique in our survey.
The remainder of the paper is organized as follows. Section 2 introduces the paper search and selection process. Section 3 first provides a workflow of deep learning-based report generation, then analyses the techniques in each component of the workflow, and finally introduces the overall training strategies. Next, in sections 4 and 5, we introduce popular public datasets and evaluation methods, including metrics and expert evaluation. Section 6 compares the model performance of several papers under the same experimental setting. Lastly, we discuss challenges and perspectives on this topic in section 7 and provide a conclusion in section 8.
@@ 2 Search and selection of articles
Three search engines (Google Scholar, PubMed, and Springer) and four queries were employed to collect articles. They included “radiology report generation", (medical OR medicine OR health OR radiology) AND (report OR description OR caption) AND generation, modal AND (medical OR medicine OR health OR radiology) AND (report OR description OR caption) AND generation, and “medical report generation". Following the searches, the titles and abstracts of each article were read briefly to identify those that met the selection criteria. If there was uncertainty, the article was included to ensure relevant studies were not omitted. The selection criteria were framed around three aspects. First, we included articles published in the years 2021, 2022, 2023, and 2024 due to the significant number of developments using multi-modal technology in recent years. We aim to focus on the latest algorithms not covered in previous surveys. Second, the studies must be original researches focused on the automatic generation of full-text natural language radiology reports and include quantitative evaluation results. Techniques generating short captions of one or two sentences are excluded due to the differing nature of long report and sentence generation. Third, papers published in journals, conferences, and conference workshop proceedings were included. Moreover, papers uploaded on the arXiv website in 2023 and 2024 with over 30 citations were also selected. A total of 144 papers were identified using three search engines. In addition, by tracing the ancestry and descendants of papers, we identified another 24 papers. After removing duplicates, 97 publications were retained. We thoroughly read these works and applied exclusion criteria. First, at least one of the generated languages should be English. Second, at least one of the input data should be images. Finally, 89 works were included in the following analysis.
@@ 3 Methods
Deep learning-based radiology report generation typically follows a standard workflow summarized in Figure 2. This section analyzes the techniques in the 89 works, based on the workflow identified. Overall, a basic radiology report generation framework consists of 5 steps: (1) multi-modality data acquisition (section 3.1); (2) data preparation (section 3.2); (3) feature learning (section 3.3); (4) feature fusion and interaction (section 3.4); and (5) report generation (Section 3.5). In addition, novel training strategies, including modifying loss functions, reinforcement learning, and curriculum learning, are described in Section 3.6.
Medical images, when analyzed with or without other types of data, are firstly prepared (step 2). Subsequently, they are input into feature extractors to perform feature learning (step 3), predominantly implemented using CNN or Transformer architectures, along with multiple enhancement modules (e.g., auxiliary task and contrastive learning). The feature extractors aim at extracting features relevant to report generation, and the enhancement modules are utilized to improve the expressiveness of the features. For certain approaches, a feature fusion and interaction module (step 4) is subsequently applied to align cross-modal data and mitigate the negative effects caused by differences between the visual and textual domains. After fusion and interaction, the features are conveyed back to the feature extractor or directly input into the generator to generate the report (step 5). The training strategy is implemented to improve learning effectiveness during training. Table LABEL:tab:overall in Appendix A presents detailed information for each paper across the five steps and training strategies.
3.1 Multimodality input data
The input data refers to the data received by the report generation system. During model training and inference, the data can vary; for example, both images and real reports are used as inputs during training, while only images are used during inference (Shetty et al., 2023). However, the model learns from the distribution and features of the training data. If the testing data changes, mostly, the model will struggle to generalize, resulting in decreased performance. Therefore, in this section, we mainly introduce the acquisition of input data that is consistent between the training and testing phases in the reviewed papers. The input data includes:
@@ 3.1 Multimodality input data
  • Medical terminology refers to medical terms and expressions, which are from keyword labels of images (Huang et al., 2021a, b, 2022; Liu et al., 2023b), or self-built corpora (Liu et al., 2021b; Cao et al., 2022, 2023; Xue et al., 2024; Gu et al., 2024; Li et al., 2023c), storing common descriptions found in medical reports.
@@
  • Medical knowledge base mostly records the connections between different organs and diseases, and is presented in a graph format. Graph is a fundamental data structure consisting of a set of nodes and edges, which can easily represent a set of subjects and their connections. The knowledge graph can primarily be obtained in two ways: 1) public datasets such as the RadGraph (Jain et al., 2021) used by two works (Yang et al., 2022; Li et al., 2023a); 2) self-designed knowledge bases according to authoritative medical standards (Huang et al., 2023; Xu et al., 2023) or disease labels (Jia et al., 2022). Yang et al. (2023) argued that manual graph construction is limited to diseases, further complicating the adaptation of these models for diverse datasets. To overcome it, they automatically constructed the medical knowledge based on real reports during training and leveraged the knowledge base during model inference.
  • Real texts report is mostly obtained by data retrieval (Liu et al., 2021b; Song et al., 2022; Yang et al., 2022; Li et al., 2023a; Liu et al., 2023b; Li et al., 2023c; Jin et al., 2024). For example, for each input image, Liu et al. (2021b) retrieved similar images from the training dataset and utilized the corresponding reports. This process mimics radiologists consulting previous medical case reports when drafting their own. In addition, Liu et al. (2023b) also obtained pre-defined sentences based on input terminologies.
  • Clinical information encompasses patient demographics (e.g., age and gender), clinical observations, and medical histories. It is included within the indication section of radiology reports (see Figure 4).
  • Questionnaires: report generation model can be trained in a visual question answering way (Tanwani et al., 2022; Pellegrini et al., 2023). The questionnaires are provided by public datasets, such as VQA-Rad (Lau et al., 2018) and Rad-ReStruct (Pellegrini et al., 2023).
@@ 3.2 Data preparation
Data preparation endeavors to enhance data quality and prepare it for model deployment, typically encompassing data cleansing, transformation, and organization. Conventional preparation methods include image resizing and cropping, text tokenizing, converting all tokens to lowercase, removing non-alphabetic tokens, and implementing data augmentation procedures. The methods utilized in each paper are outlined in Table LABEL:tab:overall in Appendix A, but it is worth noting that conventional preparation methods are so ubiquitous that some papers do not mention them. While the information is not recorded in the table, it does not mean the absence of a data preparation process.
Novel data preparation methods in the reviewed papers can be categorized into filtering (Ramesh et al., 2022) and grouping (Wang et al., 2022b). Ramesh et al. (2022) argued that writing a radiology report necessitates referencing historical information, which inevitably included descriptors such as ‘again’ and ‘decrease’. However, these terms cannot be inferred from a single image, therefore Ramesh et al. filtered such descriptions in the reports. This exclusion was found to facilitate the model’s learning process.
@@
Grouping refers to organizing sentences from ground truth reports into distinct sections, typically relying on keywords from pre-defined knowledge graphs and filtering rules. Each section describes a specific anatomical structure. Grouping aims to enable the generation system to process various types of sentences differently. For instance, Wang et al. (2022b) employed different decoders to generate descriptions for different anatomical structures. Alongside the reviewed papers, a recently released public dataset named ImaGenome (Wu et al., 2021) also includes grouping results in their annotation files (see Section 4). The grouping result becomes more easily accessible.

### v2

@@ Abstract
Automatic radiology report generation can alleviate the workload for physicians and minimize regional disparities in medical resources, therefore becoming an important topic in the medical image analysis field. It is a challenging task, as the computational model needs to mimic physicians to obtain information from multi-modal input data (i.e., medical images, clinical information, medical knowledge, etc.), and produce comprehensive and accurate reports. Recently, numerous works have emerged to address this issue using deep-learning-based methods, such as transformers, contrastive learning, and knowledge-base construction. This survey summarizes the key techniques developed in the most recent works and proposes a general workflow for deep-learning-based report generation with five main components, including multi-modality data acquisition, data preparation, feature learning, feature fusion and interaction, and report generation. The state-of-the-art methods for each of these components are highlighted. Additionally, we summarize the latest developments in large model-based methods and model explainability, along with public datasets, evaluation methods, current challenges, and future directions in this field. We have also conducted a quantitative comparison between different methods in the same experimental setting. This is the most up-to-date survey that focuses on multi-modality inputs and data fusion for radiology report generation. The aim is to provide comprehensive and rich information for researchers interested in automatic clinical report generation and medical image analysis, especially when using multimodal inputs, and to assist them in developing new algorithms to advance the field.
@@ 1 Introduction
Medical images can offer detailed information about the bodies and help physicians screen, diagnose, and monitor medical conditions without requiring invasive techniques (Beddiar et al., 2023; Liao et al., 2023). Radiologists summarize the information extracted from medical imaging to radiological reports for clinical decision-making. The manual generation of reports is however labour-intensive, time-consuming, and requires extensive expertise (Beddiar et al., 2023). The volume of radiological imaging data is growing significantly faster than the number of trained radiologists. In some cases, each radiologist needs to analyze an image every 3 to 4 seconds during an 8-hour workday (Hosny et al., 2018; McDonald et al., 2015). Rimmer (2017) revealed that 97% of the United Kingdom radiology departments faced challenges in meeting diagnostic reporting requirements. Topol (2019) pointed out that the demand for medical image interpretation greatly surpasses the current capacity of physicians in the United States. This makes it challenging for radiologists to provide high-quality reports within the scheduled time. The current demand extends patient waiting time increases the risk of disease transmission (Beddiar et al., 2023) and compromises patient care. The development of automatic report-generation techniques can help alleviate this problem.
Automatic high-quality report generation is challenging. It is intrinsically a multi-modality problem. (Tu et al., 2024; Yan et al., 2023). A high-standard report demands clarity, correctness, conciseness, completeness, consistency, and coherence (Kaur et al., 2022; Liao et al., 2023). To meet these requirements, radiologists need to combine image data with information from other modalities, such as medical knowledge and clinical history, since a single image lacks sufficient diagnostic details. For example, the same image may yield different diagnoses depending on the associated clinical context. Ideally, automated systems should mimic this process, and experiments also show that diverse modalities can enhance report generation performance (Yang et al., 2022; Li et al., 2023b; Zhang et al., 2024a).
Mimicking this process is difficult for automated systems. First, multi-modal data are represented in distinct feature spaces with different sizes, data formats (e.g. discrete vs. continuous), and data structures (e.g. graphs, vectors, etc.) (Li et al., 2023e). This disparity complicates the alignment and integration of multi-modal data. Second, models need to remove redundant information while identifying and leveraging complementary information from multi-modal inputs. Third, different modalities exhibit various types of noise. Fourth, data from disparate sources contribute unevenly to report generation, necessitating that models assess and appropriately weigh these contributions. Due to these challenges, previously developed techniques mostly considered images as input, while for the past three years, multi-modality deep learning developed very rapidly. An increasing number of research papers endeavoured to emulate physicians by leveraging multi-modal data for the generation of diagnostic reports. Additionally, the recent rise of large models that utilize prompts has made the use of multi-modal inputs a mainstream trend, as shown in Figure 1.
@@
Most of the previous surveys on this topic (Kaur et al., 2022; Beddiar et al., 2023; Liao et al., 2023; Shamshad et al., 2023; Liu et al., 2023a; Pang et al., 2023) did not include non-imaging inputs. Messina et al. (2022) considered non-imaging inputs but only included 6 papers. In total, the previous surveys included 40 to 66 papers for report generation, primarily focusing on articles published before 2022. In these studies, automated systems typically follow five key steps: multi-modality input acquisition, data preparation, feature learning, feature fusion and interaction, and report generation. Performance-improving techniques are applied at specific stages, making step-wise categorization useful for clarifying their functions. Although earlier surveys have addressed this point, they have not provided sufficient clarity. Additionally, we found that previous surveys have given limited attention to two important issues. While the application of large medical models in report generation has gained increasing recognition, current surveys lack analysis of their usage and fine-tuning. Further, interpretability and uncertainty quantification in clinical decision making remains a concern. While Messina et al. (2022) study on explainable AI is comprehensive, there is limited survey analysis of papers published after 2022.
This survey differs from previous ones in five main contributions:
  • We analyse an additional 22 papers that utilize non-image inputs and focus on the acquisition, analysis, and integration of multi-modal inputs. To the best of our knowledge, it is the first review to investigate state-of-the-art multi-modal data processing techniques for report generation.
  • Totally, we examine 100 papers published from 2021 to 2024 to provide a comprehensive study on novel techniques in automatic report generation, of which 90 focus on traditional deep learning methods, and 10 are based on large model approaches.
  • We propose a general workflow for deep-learning-based report generation with a taxonomy of approaches employed, as shown in Figure 2. Table LABEL:tab:overall in Appendix A summarizes all reviewed papers.
  • For large model-based approaches, including language models, vision models, and multimodal models, we summarize the acquisition of image-text datasets with prompts, the evolution from large language model to large multimodal model, and how to train large medical models with limited resources.
  • We present recent advancements in explainability for report generation, including visualization, auxiliary tasks, and uncertainty quantification.
Figure 2: An overview of report generation: workflow and taxonomy of employed approaches. a. The typical workflow of automatic radiology report generation. b. The summary of the usage of techniques in the reviewed papers for each step. (x%) represents the percentage of these articles relative to the 100 papers reviewed.
The remainder of the paper is organized as follows. Section 2 introduces the paper search and selection process. Section 3 first provides a workflow of report generation, then analyses the techniques in each component of the workflow. Section 4 focuses on the application of large models in report generation. Section 5 discusses the latest developments in explainability. Next, in sections 6 and 7, we introduce popular public report generation datasets and evaluation methods, including metrics and expert evaluation. Section 8 compares the model performance of several papers in the same experimental setting. Lastly, we discuss challenges and perspectives on this topic in Section 9 and provide our conclusions in Section 10.
@@ 2 Search and selection of articles
Three search engines (Google Scholar, PubMed, and Springer) and four queries were employed to collect articles. They included “radiology report generation", (medical OR medicine OR health OR radiology) AND (report OR description OR caption) AND generation, modal AND (medical OR medicine OR health OR radiology) AND (report OR description OR caption) AND generation, and “medical report generation". Following the searches, the titles and abstracts of each article were read briefly to identify those that met the selection criteria. If there was uncertainty, the article was included to ensure relevant studies were not omitted. The selection criteria were framed around three aspects. First, we included articles published in the years 2021, 2022, 2023, and 2024 due to the significant number of developments using multi-modal technology in recent years. We aim to focus on the latest algorithms not covered in previous surveys. Second, the studies should be original research on the automatic generation of full-text natural language radiology reports. While large model-based methods may generate only parts of reports for specific prompts, they are included in this survey because they have the potential to produce complete reports with different prompts. In contrast, methods that generate short captions of one or two sentences are excluded due to the differing nature of long reports and sentence generation. Third, papers published in journals, conferences, and conference workshop proceedings were included. Additionally, papers from arXiv in 2023 and 2024 with more than 15 citations were selected. Due to the potential impact of emerging medical large models on report generation, relevant articles from arXiv are included regardless of their citation count.
A total of 151 papers were identified using three search engines. In addition, by tracing the ancestry and descendants of papers, we identified another 30 papers. After removing duplicates, 110 publications were retained. We thoroughly read these works and applied exclusion criteria. First, at least one of the generated languages should be English because it is the common language of academia, which ensures broader accessibility, easier evaluation, and more accurate comparisons of algorithm performance. Second, at least one of the input data should be images. Finally, 102 works were included in the following analysis.
@@ 3 Methods
Deep-learning-based radiology report generation typically follows a standard workflow summarized in Figure 2 a. Overall, a basic radiology report generation framework consists of 5 steps: (1) multi-modality data acquisition (Section 3.1); (2) data preparation (Section 3.2); (3) feature learning (Section 3.3); (4) feature fusion and interaction (Section 3.4); and (5) report generation (Section 3.5). Although traditional deep learning methods and large model-based approaches prioritize different aspects of research for automatic report generation, both adhere to the workflow shown in Figure 2 a. Traditional deep learning methods focus on enhancing the effectiveness of each step, while the exploration of large models is in its early stages. Large models focus on large-scale data acquisition, the transition from large language models to large multimodal models, and effective training. This section analyzes the techniques in the 100 works based on the workflow, and the next section discusses the details of large models. We believe that effective techniques from traditional methods can inspire the future development of large models due to their similar workflows.
Typically, medical images, when analyzed with or without other types of data, are first prepared, including image normalization, resizing, etc. (step 2). Subsequently, they are inputted into feature extractors to perform feature learning (step 3), predominantly implemented using a Convolutional Neural Network (CNN) or Transformer architectures, along with auxiliary loss functions or enhancement modules. The feature extractors aim at extracting features relevant to report generation, and additional techniques are utilized to improve the expressiveness of the features. For certain approaches, a feature fusion and interaction module (step 4) is subsequently applied to align cross-modal data and mitigate the negative effects caused by differences between the visual and textual domains. After fusion and interaction, the features are input into the generator to generate the report (step 5). Throughout this process, the network parameters are generally trainable. In some cases, a frozen pre-trained encoder can be used to enhance training stability. For instance, the visual encoder in large-model-based methods (Li et al., 2024; Lee et al., 2023). Table LABEL:tab:overall in Appendix A presents detailed information for each paper across the five steps and training strategies.
3.1 Multi-modality inputs acquisition
The input data refers to the data received by the report generation system. During model training and inference, the data can vary; for example, both images and real reports are used as inputs during training, while only images are used during inference (Shetty et al., 2023). However, the model learns from the distribution and features of the training data. If the testing data changes the model will likely struggle to generalize, resulting in decreased performance. Therefore, in this section, we define and explain the acquisition of input data used consistently in both the training and testing phases of the reviewed papers.
Input data includes images, medical terminology, medical knowledge base, real text reports, clinical information, questionnaires, and medical descriptions. Figure 3 illustrates the types, examples and sources of these data. The following provides a detailed description:
Figure 3: A summary of the types, examples, and sources of multi-modal input data.
@@ 3.1 Multi-modality inputs acquisition
  • Instructions refer to the explanations provided for completing a task or carrying out an operation. It comprises prompts and questionnaires. Prompts like “Generate a comprehensive and detailed diagnosis report for this chest x-ray image” are commonly used with large models for report generation (Tanwani et al., 2022; Pellegrini et al., 2023b; Lee et al., 2023; Wang et al., 2023e; Hyland et al., 2023; Bannur et al., 2024; Chen et al., 2024; Guo et al., 2024; Zhang et al., 2024a; Li et al., 2024; Alkhaldi et al., 2024). Questionnaires are sets of questions designed for report generation (Tanwani et al., 2022; Pellegrini et al., 2023a). Pellegrini et al. (2023a) developed a structured report dataset for Chest X-rays, containing hundreds of questions, in which reports are generated by answering a series of questions, similar to the process of hierarchical visual question answering.
@@
  • Medical terminology refers to medical terms and expressions, which are from keyword labels of images (Huang et al., 2021a, b, 2022; Liu et al., 2023c), or self-built corpora (Liu et al., 2021b; Cao et al., 2022, 2023; Xue et al., 2024; Gu et al., 2024; Li et al., 2023d), storing common descriptions found in medical reports. Incorporating medical terminologies can assist the model in generating descriptions that are relevant to medical terminology. Liu et al. (2023c) further developed a concepts library for medical descriptions based on various terminologies, providing detailed information on symptoms and health status. For example, the term “Cardiomegaly” corresponds to “The heart size is enlarged”.
  • Real texts report refers to radiology reports written by physicians. In the reviewed papers, two methods are used to obtain these reports. The first retrieves them from the training dataset to avoid overlap with the test set, as detailed in Section 3.2. It mimics the behaviour of radiologists referencing similar prior cases when writing reports. The second uses prior reports, referring to the patient’s historical data (Bannur et al., 2024). As described in Section 3.2, utilizing this data enables monitoring of temporal changes, which is essential for clinical applications.
@@
  • Medical knowledge base mostly records the connections between different organs and diseases, and is presented in a graph format. A graph is a fundamental data structure consisting of a set of nodes and edges, which can easily represent a set of subjects and their connections. A knowledge graph can primarily be obtained in two ways: (1) public datasets such as the RadGraph (Jain et al., 2021) used by two works (Yang et al., 2022; Li et al., 2023b). Recently, an enhanced version of RadGraph, termed RadGraph-XL (Delbrouck et al., 2024), was introduced, offering improved performance. This update extends its applicability to a broader range of medical imaging modalities (i.e., chest CT, abdomen/pelvis CT, brain MRI, and chest X-ray), anatomical regions, and data sources; (2) self-designed knowledge bases according to authoritative medical standards (Huang et al., 2023; Xu et al., 2023) or real reports in training dataset (Jia et al., 2022). Utilizing a knowledge base allows the model to learn medical knowledge akin to a physician.
  • Clinical information encompasses patient demographics (e.g., age and gender), clinical observations, and medical histories. By incorporating clinical information, the model can obtain rich clinical context, which helps it better understand the specific situation and background of the patient. Clinical information is embedded in the indication, technique, and comparison sections of radiology reports (see Figure 7). While most studies utilize only the indication section (Zhou et al., 2021; Dalla Serra et al., 2022, 2023a, 2023b; Hyland et al., 2023), MAIRA-2 (Bannur et al., 2024) leverages all three sections, stating that the technique and comparison sections provide additional information, and demonstrates their positive impact through experiments.
@@ 3.2 Data preparation
Data preparation endeavours to enhance data quality and prepare it for model deployment, typically encompassing data cleansing, transformation, and organization. Conventional data preparation techniques primarily target images and text. Image processing typically involves resizing and cropping, while text processing includes case normalization, removal of non-alphabetic tokens, and tokenization. The methods utilized in each paper are outlined in Table LABEL:tab:overall in Appendix A, but it is worth noting that conventional preparation methods are so ubiquitous that some papers do not mention them. While the information is not recorded in the table, it does not mean the absence of a data preparation process.
Novel data preparation methods in the reviewed papers can be categorized into filtering (Ramesh et al., 2022), grouping (Wang et al., 2022b), and case-specific information extraction. Ramesh et al. (2022) argued that writing a radiology report necessitates referencing historical information, which inevitably included descriptors such as ‘again’ and ‘decrease’. However, these terms cannot be inferred from a single image, therefore Ramesh et al. (2022) filtered such descriptions in the reports. This exclusion was found to facilitate the model’s learning process; however, simply modifying ground truth reports does not accurately reflect clinical practice. In some clinical settings, monitoring temporal changes is essential for the effective diagnosis and management of various diseases, such as pneumonia (Kalil et al., 2016; Wu et al., 2021a). At the same time, follow-up reports are often shorter than those for the initial image, focusing only on changes or new findings (e.g., “Otherwise, little change.”). Without historical context can result in incomplete image descriptions. Current publicly available datasets are typically organized by subject identifiers and study identifiers, representing different examinations for the same patient and arranging them chronologically (Johnson et al., 2019a; Wu et al., 2021a; Chambon et al., 2024). Analyzing a sequence of images or inputting prior reports into the model might be more appropriate (Bannur et al., 2024; Dalla Serra et al., 2023a). Another approach is to train using only the initial image-report pairs to maintain data validity, though this would limit the model’s applicability and significantly reduce the amount of available training data.
@@
Grouping refers to organizing sentences from ground truth reports into distinct sections, typically relying on keywords from pre-defined knowledge graphs and filtering rules. Each section describes a specific anatomical structure. Grouping aims to enable the generation system to process various types of sentences differently. For instance, Wang et al. (2022b) employed different decoders to generate descriptions for different anatomical structures. Alongside the reviewed papers, a recently released public dataset named ImaGenome (Wu et al., 2021a) also includes grouping results in their annotation files (see Section 6). The grouping result becomes more easily accessible.
Figure 4: Three typical methods for extracting case-related information. (a) Extracting relevant textual reports from the training set by comparing the similarity between the input image and images in the training set. (b) Extracting relevant medical descriptions from a concepts library based on the image’s annotated keywords. (c) Automatically extracting entities (e.g., anatomy) from the reports retrieved in (a) and using these entities to search the knowledge base for related knowledge graphs. “…” refers to concepts such as these. The figure only lists a few examples from the training dataset, keyword labels of images, medical descriptions, recognized entities in the report, and knowledge base.
When using medical knowledge bases and real text reports, inputting the entire knowledge base or all reports from the training dataset often leads to information redundancy, making it difficult for the network to extract relevant information. Extracting case-specific information can alleviate this problem. Figure 4 illustrates three highly effective methods in reviewed papers.
The first method is the retrieval of related real text reports (Liu et al., 2021b; Song et al., 2022; Yang et al., 2022; Li et al., 2023b; Liu et al., 2023c; Li et al., 2023d; Jin et al., 2024). This process mimics radiologists consulting previous medical case reports when drafting their own. By calculating the similarity between the input image and the images in the training set, we select the reports corresponding to the top $N$ most similar images and use these reports as inputs. The second method is the retrieval of related medical descriptions. Liu et al. (2023c) obtained medical description from a pre-defined concepts library based on input terminologies. However, this method still depends on disease tags from radiologists for high performance, increasing their workload.
The third method combines retrieved reports in the first method with a medical knowledge base (e.g., RadGraph) to extract case-related medical knowledge (Yang et al., 2022; Li et al., 2023b). An automated tool first extracts entities from the reports, and then all related information from the knowledge base is retrieved as input. Alternatively, Li et al. (2023d) use the annotation tool in RadGraph to extract entities and positional information from retrieved reports.
Additionally, two studies (Xu et al., 2023; Huang et al., 2023) used image classification results to process self-built graphs and extract case-related information. This method is less effective than the three mentioned above, partly due to inherent classification errors. Improving classifier performance could enhance the quality of report generation.